OpenStudy (anonymous):

Write the other three proportions using the four values in each proportion: a). 2/5 = 10/25 b). a/9 = 12/27

OpenStudy (anonymous):

a. 5/2 = 25/10 2/10 = 5/25 10/2 = 25/5 b. 9/a = 27/12 a/12 = 9/27 12/a = 27/9 a = 4